At auction, 8-month T-bills were sold at a discount of 7.410%. What was the simple annual yield r? HINT [See Example 6.] (Round your answer to three decimal places.)
Answer #1

Answer is 11.115%

let us assume the par value as 1000

purchase price = (par value - (par value * discount rate)) = (1000-(1000*7.410%) = 925.90

simple annual yield = {(par value - purchase price) / par value } * (360 *100/ days to maturity)

= {(1000 - 925.9) / 1000 } * (360 *100/ 240)

= {(74.10) / 1000 } * (360 *100/ 240)

= 0.0741 * (150)

= 11.115 %

Thank You